### v1

# Designing Elegant Bell Inequalities

Abstract
Elegant Bell inequality is well known for its much exploited property, being maximally violated by maximal entanglement, mutually unbiased bases, and symmetric informationally complete positive operator-valued measure elements. It is the only one with such property known so far. We present a method to construct Bell inequalities with violation feature analogous to original elegant Bell inequality in high dimension from a simple analytic quantum bound. A Bell inequality with such feature is derived in three dimension for the first time. It shows larger violation than existing Bell inequalities of similar classes while requiring arguably small number of measurements.
I Introduction
Performing local measurements on the shared system, distant subsystems can reveal a quantum correlation that cannot be reproduced by any local hidden variable theoretical explanation. Such nonlocality of the correlation can be tested via the violation of Bell inequalities [1, 2]. What Bell inequality can tell us about the system goes beyond the scope of the mere existence of nonlocal correlation. It has been investigated the method to exploit Bell’s theorem for further certifying the properties of given system in so-called device-independent (DI) approach [3, 4]. This approach leads to various applications, including DI quantum key distribution [4, 5, 6, 7, 8] and DI certification of randomness [9, 10, 11].
Reflecting on the aforementioned possibilities provided by Bell inequalities, it is of prime importance to derive them such that their maximal violations are obtained from desirable features, for example maximal entanglement [12, 13, 14, 15, 16, 17, 18, 19, 20], mutually unbiased bases (MUBs) [19, 20], or symmetric informationally complete POVM elements (SICs) [21, 20]. As is well known, Clauser-Horne-Shimony-Holt Bell inequality is the case in which the maximal violation can be obtained from maximal entanglement and Pauli operators having MUBs as their eigenstates. There exists another bipartite Bell inequality so-called Gisin’s elegant Bell inequality (EBI) [21],
$E_{11}+E_{12}-E_{13}-E_{14}$
$+E_{21}-E_{22}+E_{23}-E_{24}$
$+E_{31}-E_{32}-E_{33}+E_{34}\leq 6,$
also optimally violated by maximally entangled state. The expectation value of the product outcomes of Alice’s $x$-th measurement and Bob’s $y$-th measurement is denoted by $E_{xy}$. Quantum bound for EBI is proven as $64\sqrt{3}\simeq 6.9282>$ 6 [22]. What makes EBI more interesting is that its violation is obtained from the case where the local measurements of each subsystem is constructed by MUBs and SICs respectively. As first remarked by Gisin, the geometrical symmetries in the optimal measurement settings of EBI are clearly revealed when they are mapped on Bloch’s sphere [21]. Three eigenstates defining three projective measurements of Alice form three mutually orthogonal vectors on Bloch sphere and four eigenstates of Bob’s side make a tetrahedron [21]. Original EBI has been used in several fields of quantum information, including DI randomness certification [22, 23], DI certification of non-projective measurement [24, 25], and for other tasks [26, 27, 28]. Self-testing property of EBI has also been studied in [29].
MUBs and SICs are important resources in quantum cryptography [30, 31, 32, 22, 23]. There has been efforts to derive Bell inequalities for MUBs or SICs exploiting the advantage of high dimension [19, 20]. For DI randomness certification, for which high dimension provides natural advantage [22], Bell inequalities for SICs has been used [22, 23, 20]. EBI is distinguished from ones either for MUBs or SICs as it is optimally violated by both of them.
We address the problem of deriving Bell inequalities with violation feature analogous to original EBI in high dimensions. The problem is closely related to the question of what benefit for Bell nonlocality would come from simultaneous consideration of the symmetries of MUBs and SICs. It is worth noting that although EBI is originally derived in the different context [33, 21], we focus on the generalization of the property, being maximally violated by maximal entanglement, MUBs and SICs. Such generalization of EBI has not much been investigated so far to our knowledge. We approach the raised problem by defining a generalized type of correlation as follows.
Definition. For prime $d\geq 3$, generalized EBI correlation is defined as the bipartite correlation obtained from the quantum realization: (i) maximally entangled state $\left|\phi_{d}^{+}\right>:=(1/\sqrt{d})\sum_{\alpha=0}^{d-1}\left|\alpha\alpha\right>$, (ii) Alice’s $d^{2}-1$ observables comprising $\mathcal{W}_{d}\setminus\mathds{1}$ and (iii) Bob’s $d^{2}$ observables having $d^{2}$ SICs in their eigenspaces.
In condition (ii), $\mathcal{W}_{d}:=\{X^{m}Z^{n}|m,n\in[0,d)\}$ is the set of $d^{2}$ Weyl-Heisenberg (WH) operators defined with $X:=\sum_{\alpha=0}^{d-1}\left|\alpha+1\right>\left<\alpha\right|$, $Z:=\sum_{\alpha=0}^{d-1}\omega^{\alpha}\left|\alpha\right>\left<\alpha\right|$ and $\omega:=e^{2\pi i/d}$. The integer domain $\{0,1,\ldots,d-1\}$ is denoted by $[0,d)$ and similar abbreviation is to be used throughout this work. Note that $X^{0}Z^{0}=\mathds{1}$ is not included in the condition (ii). Conditions (i) and (iii) are natural generalization of optimal quantum realization of original EBI. As for the generalization of Alice’s optimal observables, two interpretations are possible. The set of Alice’s optimal observables for EBI, $\sigma_{x},\sigma_{y},\sigma_{z}$, [22] can be viewed either as “$d+1$ observables” defining $d+1$ sets of MUBs, or as “$d^{2}-1$ unitary operator bases” [34] for a traceless operator $\vec{a}\cdot\vec{\sigma}$ defined with a three-dimensional unit vector $\vec{a}$ and $\vec{\sigma}:=(\sigma_{x},\sigma_{y},\sigma_{z})$ when $d=2$. The above duality arises as $d+1=d^{2}-1$ when $d=2$. Condition (ii) clearly generalizes the latter. Under condition (ii), complete MUBs are realized in eigenspace of $d+1$ among $d^{2}-1$ observables of Alice, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [35]. Moreover, the eigenbases of all $d^{2}-1$ observables of Alice comprises the $d-1$ same sets of complete $d+1$ MUBs. It can be easily seen from the fact that $XZ^{n}$ and $X^{m}Z^{mn}$ has the same eigenbases for all $m$ and $\mathcal{W}=\{X^{m}Z^{mn}|m,n\in[0,d)\}$ for prime $d$. In the following, we will further explain the reason for the choice of optimal observables of Alice as unitary operator bases.
Figure 1: On the shared system, two subsystems, Alice and Bob, respectively performs one of $d^{2}-1$ and $d^{2}$ measurements outputting one of $d$ outcomes.
Another insight that underlies our approach is to exploit $systematic$ and $versatile$ features of WH group covariant SICs. For brevity, consider the parameterization $W_{dp+q}:=X^{p}Z^{q}$, such that $\mathcal{W}_{d}=\{W_{j}|j\in[0,d^{2})\}$. Group covariant SICs is derived by acting group elements on a single state, so-called fiducial vector [36]. For the case of WH group covariant SICs, group elements $W_{j}$ generates $d^{2}$ SICs on the support of fiducial state $\left|\varphi\right>$, $\{W_{j}\left|\varphi\right>|j\in[0,d^{2})\}$. Weyl-Heisenberg group covariance plays crucial role for satisfying the condition (iii) and provides systematic feature for our formalism. Versatility of our method is from the fact that it can be applied only if WH group covariant SICs for given local dimension exists. The existence of SICs are known in all dimension up to $d=193$ [37]. More importantly, all group covariant SICs is covariant to WH group in prime dimensions [38].
II Preliminaries
We consider the measurement scenario of two subsystems Alice and Bob respectively performing one of possible $d^{2}-1$ and $d^{2}$ number of $d$ outcome measurements $A_{x}$ and $B_{y}$ on the shared system. Measurements are labelled with $x\in\{1,2,\ldots,d^{2}-1\}$ and $y\in\{0,1,\ldots,d^{2}-1\}$. Outcomes of $A_{x}$ and $B_{y}$ are respectively distinguished with indices $\alpha,\beta\in\{0,1,\ldots,d-1\}$. The scenario is illustrated in Fig. 1. Probability of obtaining the outcomes $\alpha$ and $\beta$ respectively from settings $x$ and $y$ is denoted by $P(\alpha\beta|xy)$. A statistics of the Bell test can be described with correlation, $\mathbf{p}:=\{P(\alpha\beta|xy)|\forall\alpha,\beta,x,y\}$. Quantum correlation allows
$P(\alpha\beta|xy)=\operatorname{tr}\left[\rho_{AB}(G_{\alpha}^{x}\otimes H_{\beta}^{y})\right]$
where $\rho_{AB}$ is the bipartite density matrix of Alice and Bob. Measurement operators, $\{G_{\alpha}^{x}\}$ and $\{H_{\beta}^{y}\}$, define the measurements of Alice and Bob respectively. Bell inequality is used to show that there is quantum correlation which does not allow the local hidden variable (LHV) theoretical description of the correlation [1]. Geometrically, LHV correlation region forms a polytope, $\mathcal{L}$, and quantum region, $\mathcal{Q}$, is known to include it [2].
The generic form of so-called Bell expression [2] without marginal probabilities can be defined as,
$S$ $=\sum_{\alpha,\beta}\sum_{x,y}g_{x,y}^{\alpha,\beta}P(\alpha\beta|xy)$ (1)
$=\sum_{n}\sum_{x,y}f_{x,y}^{n}\langle A_{x}^{n}B_{y}^{n}\rangle.$ (2)
where $n\in(1,d)$ and $g_{x,y}^{\alpha,\beta}\in\mathbb{R}$ is the real weight for each probability. For simplicity, the summations over all possible values are to be denoted without limits. The correlator $\langle A_{x}^{n}B_{y}^{n}\rangle=\sum_{\alpha,\beta=0}^{d-1}\omega^{n(\alpha+\beta)}P(\alpha\beta|xy)$, where the outcomes of the measurements $A_{x}$ and $B_{y}$ are set as $\omega^{\alpha},\omega^{\beta}\in\{1,\omega,\omega^{2},\ldots,\omega^{d-1}\}$. And the coefficient $g_{x,y}^{\alpha,\beta}:=\sum_{n}f_{x,y}^{n}\omega^{n(\alpha+\beta)}$ The condition $f_{x,y}^{d-n}:=(f_{x,y}^{n})^{*}$ guarantees the real-valuedness of $S$. A expression is specified by a $(d^{2}-1)\times d^{2}$ coefficient matrices, $F_{n}$, whose entries are $\{f_{x,y}^{n}\}$. We define $x$-th row and $y$-th column of $F_{n}$ respectively as $\mathbf{r}_{x}^{n}:=(f_{x,0}^{n},f_{x,1}^{n},\ldots,f_{x,d^{2}-1}^{n})^{\intercal}$ and $\mathbf{c}_{y}^{n}:=(f_{1,y}^{n},f_{2,y}^{n},\ldots,f_{d^{2}-1,y}^{n})^{\intercal}$ such that $F_{n}=[\mathbf{r}_{1}^{n},\mathbf{r}_{2}^{n},\ldots,\mathbf{r}_{d^{2}-1}^{n}]^{\intercal}=[\mathbf{c}_{0}^{n},\mathbf{c}_{1}^{n},\ldots,\mathbf{c}_{d^{2}-1}^{n}]$. In terms of $F_{n}$, real-valuedness condition is re-written as $F_{n}=F_{d-n}^{*}$.
Maximal value of $S$ under LHV theory is $L:=\max_{\mathbf{p}\in\mathcal{L}}S(\mathbf{p})$. Violation of the Bell inequality by quantum mechanics, $L<Q$, is shown with the maximal quantum value, $Q:=\max_{\mathbf{p}\in\mathcal{Q}}S(\mathbf{p})$. Quantum mechanical expectation can be evaluated with the operator,
$\mathcal{B}=\sum_{n}\sum_{x,y}f_{x,y}^{n}A_{x}^{n}\otimes B_{y}^{n}.$ (3)
Here, we consider the quantum measurements $A_{x}=\sum_{\alpha}\omega^{\alpha}G_{\alpha}^{x}$ and $B_{y}=\sum_{\beta}\omega^{\beta}H_{\beta}^{y}$ where $G_{\alpha}^{x}$, $H_{\beta}^{y}$ are rank-one projectors.
III Main result
We outline the method to derive the condition on the correlation matrix, $F_{n}$, such that maximal quantum value is obtained by generalized EBI correlation. SOS decomposition technique is employed for analytic derivation of maximal quantum value. Consider the shifted operator, ¯$\bar{\mathcal{B}}:=Q\mathds{I}-\mathcal{B}$ and the operators $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$ defined with a real constant $Q\in\mathbb{R}$. The matrix $\mathds{I}:=\mathds{1}\otimes\mathds{1}$ is the $d^{2}\times d^{2}$ identity matrix, and $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$. The SOS decomposition ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ directly implies $\mathcal{B}\right>\leq Q$. Substituting $P_{y}^{n}$ into $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$ derives
$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}.$ (4)
In the derivation of (4), it is used the Hermiticity of $\mathcal{B}$, i.e. $\mathcal{B}=\sum_{n,y}D_{y}^{n}\otimes B_{y}^{n}=\sum_{n,y}(D_{y}^{n})^{\dagger}\otimes(B_{y}^{n})^{\dagger}$. The decomposition condition, ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$, is obtained by deriving the RHS of (4) as $Q\mathds{I}$. Saturation condition is $\operatorname{tr}[\bar{\mathcal{B}}\rho]=0$ with any valid quantum realization. It implies that $Q$ is attainable upper bound. For all $y,n$, $P_{y}^{n}\rho=0$ is equivalent to the saturation condition [19]. Importantly, this condition is satisfied by the relation $D_{y}^{n}=(B_{y}^{n})^{*}$ on the support of the maximally entangled state as $P_{y}^{n}\left|\phi_{d}^{+}\right>=[D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}]\left|\phi_{d}^{+}\right>=0$. Lemma $M\otimes\mathds{1}\left|\phi_{d}^{+}\right>=\mathds{1}\otimes M^{\intercal}\left|\phi_{d}^{+}\right>$ is used in the second equation.
Our strategy is to solve the decomposition and saturation condition by varying $f_{x,y}^{n}$ with pre-assumed generalized EBI correlation. To this end, we define the optimal observables satisfying conditions (ii) and (iii),
$A_{dp+q}$ $=\tau W_{dp+q}\quad\,\forall p,q$ (5)
$B_{y}$ $=W_{y}B_{0}W_{y}^{\dagger}\quad\forall y$ (6)
where $\tau:=\omega^{pq\delta(d,2)/2}$ [39] defined with Kronecker delta $\delta$ and the parameterization $x=dp+q$ is used in Eq.(5). The index $x\in(0,d^{2})$ is now expressed with $p,q\in[0,d)$ satisfying $(p,q)\neq(0,0)$ and $y\in[0,d^{2})$. Alice’s observables in (5) are given as $d^{2}-1$ WH operators. Bob’s observables are generated from $B_{0}$ as Eq.(6). An important part of our method is to construct $B_{0}$ such that it has the fiducial state as its eigenstate. This choice of $B_{0}$ makes it possible to realize WH group covariant SICs on Bob’s eigenspace. Consider $B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ where the eigenspace is defined as $\Phi(\varphi):=\{\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$, a set of $d$ orthonormal states including the fiducial state $\left|\varphi\right>$. Without the loss of generality, one can let $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$. Then we have,
$B_{y}$ $=\sum_{\beta}\omega^{\beta}W_{y}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|W_{y}^{\dagger}$ (7)
where the eigenbases of $B_{y}$ are now $\{W_{y}\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$. Consequently, the set of $(d-1)$-th eigenstates of all the Bob’s observables comprises the $d^{2}$ WH group covariant SICs, $\{W_{y}\left|\varphi\right>|y\in[0,d^{2})\}$.
[…]
$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (8)
when $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
Second, we solve the saturation condition with optimal realization, $\left|\phi_{d}^{+}\right>$ and observables (5) and (7). Saturation condition is given as $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}W_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=\tau^{n}(W_{1}^{n},W_{2}^{n},\ldots,W_{d^{2}-1}^{n})^{\intercal}$. The solution $\mathbf{c}_{y}^{n}$ completely derives the coefficient matrices we after. In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$. The solution is derived in the below theorem.
Theorem. Consider prime $d$, for which WH group covariant SICs is defined. Then, there always exists the expression $S$ defined with,
$f_{dp+q,dr+s}^{n}=\omega^{-n(ps+qr)}f_{dp+q,0}^{n}$ (9)
for $p,q,r,s\in[0,d)$, $dp+q\neq 0$ that satisfies $Q=\tilde{Q}$ with generalized EBI correlation.
Proof.—Consider optimal $B_{0}=\sum_{\beta=0}^{d-1}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ defined with eigenspace $\Phi(\varphi)$ containing $\left|\varphi\right>_{d-1}:=\left|\varphi\right>$. This choice of $B_{0}$ realizes SICs in the eigenspace as explained. Saturation condition is re-written with (6) as
$(B_{y}^{n})^{*}$ $=(W_{y}B_{0}^{n}W_{y}^{\dagger})^{*}$
$=\mathbf{c}_{0}^{n}\cdot(W_{y}^{*}\mathbf{A}_{n}W_{y}^{\intercal})$
$=\sum_{p,q}\omega^{-n(ps+qr)}f_{dp+q,0}^{n}(\mathbf{A}_{n})_{dp+q}$
$=\sum_{p,q}f_{dp+q,dr+s}^{n}(\mathbf{A}_{n})_{dp+q}$ (10)
where the notation $(\cdot)_{x}$ stands for the $x$-th element. The second equation is derived from saturation condition, $B_{0}^{n}=(\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n})^{*}$. Substituting $(\mathbf{A}_{n})_{dp+q}=(\tau X^{p}Z^{q})^{n}$ and $W_{y}=W_{dr+s}=X^{r}Z^{s}$ into RHS of the second equation and suitably applying the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, derives the third equation. We used the parameterizations $x=dp+q$ and $y=dr+s$ for the substitution. The equation (10) is obtained from $(B_{dr+s}^{n})^{*}=\mathbf{c}_{dr+s}^{n}\cdot\mathbf{A}_{n}$ and it results in Eq.(9).
Coefficients in Eq.(9) also satisfies the decomposition condition. Substituting (9) into $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$. Finally, showing the saturation condition is always solvable completes the theorem. The elements of $\mathbf{A}_{n}$ is the unitary operator bases for traceless operator [34]. It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$. $\blacksquare$
The above theorem states that the correlation function maximized by generalized EBI correlation always can be found from Eq.(9) with WH group covariant SICs. One has freedom in the choice of SICs, thus $B_{0}$, which determines the coefficients $\{f_{x,0}^{n}\}$ with $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$. One can construct different correlation functions by varying $\{f_{x,0}^{n}\}$ of Eq.(9) with different choices of SICs. We also note that establishing $\{A_{x}\}$ as unitary operator bases [34] for traceless operator is advantageous for restricting the number of measurements, as the bases are minimum number of unitary operators that express, in linear combination form, an arbitrary traceless unitary operator, in our case $B_{0}^{n}$.
IV Derivation of Bell inequality
IV.1 Gisin’s elegant Bell inequality
[…]
IV.2 Bell inequality for two-qutrit system
We derive a Bell inequality maximally violated by generalized EBI correlation in $d=3$. For $d=3$, there are infinitely many fiducial states generating WH covariant SICs, for example $\left|\psi\right>=1/\sqrt{2}(0,1,-\omega^{\theta})$ for $20\leq\theta\leq 1/$ 2 [41]. We choose the one with $\theta=0$, $\left|\varphi\right>:=1/\sqrt{2}(0,1,-1)=1/\sqrt{2}\left(\left|1\right>-\left|2\right>\right)$ which is known to generate the Hesse SICs [42]. With the set of orthonormal bases $\left|\varphi_{0}\right>:=\left|0\right>$, $\left|\varphi_{1}\right>:=1/\sqrt{2}(\left|1\right>+\left|2\right>)$, and $\left|\varphi_{2}\right>:=\left|\varphi\right>$, optimal $B_{0}$ is
$B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|=\begin{pmatrix}1&0&0\\ 0&-1/2&i\sqrt{3}/2\\ 0&i\sqrt{3}/2&-1/2\end{pmatrix}.$ (11)
It is worth mentioning here that our choice of $\left|\varphi_{1}\right>$ is also a fiducial state generating WH covariant SICs. By solving the saturation condition of SOS decomposition, $(B_{0})^{*}=\mathbf{c}_{0}^{1}\cdot\mathbf{A}_{1}$, the first column of $F_{1}$ is easily obtained as $\mathbf{c}_{0}^{1}=(f_{1,0}^{1},f_{2,0}^{1},\ldots,f_{8,0}^{1})^{\intercal}=(1/2)(1,1,\lambda,\nu,\mu,\lambda,\mu,\nu)^{\intercal}$ with $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. Substituting $\mathbf{c}_{0}$ into Eq.(9) and using the real-valuedness condition $F_{2}=F_{1}^{*}$, the coefficients are fully generated as in [40]. The derived coefficients can be further simplified in a functional form, $f_{dp+q,0}^{n}=1/2$ for $p=0$ and $f_{dp+q,0}^{n}=\omega^{-n(pq+3/2)+3/4}/(2\sqrt{3})$ for $p\neq 0$.
Substituting the derived coefficients into the expression (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
$S_{3}:=\sum_{n=1}^{2}\sum_{\begin{subarray}{c}p,q,r,s=0\\ 3p+q\neq 0\end{subarray}}^{2}\omega^{-n(ps+qr)}f_{3p+q,0}^{n}\left\langle A_{3p+q}^{n}B_{3r+s}^{n}\right\rangle\leq 15.$ (12)
The upper bound $L=15$ can be obtained in the set of local deterministic correlations satisfying $P(\alpha\beta|xy)=1$ for certain combination of $\alpha,\beta$ and otherwise the probability is 0. The optimal strategy is found as, for all $x$, assigning $P(00|xy)$, $P(01|xy)$, $P(02|xy)$ as 1 respectively for $y=0,1,2$, $y=3,6$, and $y=4,5,7,8$. Violation is obtained from the maximal quantum value, $\tilde{Q}=18$, evaluated from (8) with $\lVert F_{n}\rVert=9$ for $n=1,2$. The upper bound is attained by two parties performing measurements, $A_{x}=W_{x}$ and $B_{y}=W_{y}B_{0}W_{y}^{\dagger}$, on the two-qutrit maximally entangled state when $x\in(0,9)$ and $y\in[0,9)$. As expected, $2d^{2}$ eigenstates, $W_{y}\left|\varphi_{1}\right>$ and $W_{y}\left|\varphi_{2}\right>$ for all $y$, comprise dual sets of $d^{2}$ SICs and eigenstates of four optimal observables for $x=1,3,4,5$, which are given as $Z,X,XZ,XZ^{2}$, defines the complete sets of MUBs [40].
We remark that the Bell inequality defined in (12) shows stronger violation than known Bell inequalities of similar classes. The critical visibility $\nu^{c}$, the smallest visibility $\nu$ of the isotropic state $\rho=\nu\left|\phi_{d}^{+}\right>\left<\phi_{d}^{+}\right|+(1-\nu)\mathds{I}/d^{2}$ to show the violation, is 83.33% for $S_{3}$. It is smaller than that of Bell inequalities including ones defined in $d=3$ for MUBs, 96.77% [19], 96.63% [20], for SICs, 96.41% [20], and original EBI, 86.6%. In addition, the number of settings required for Alice is 8 for $S_{3}$ and it is smaller than the case of known Bell inequality for $d=3$ SICs which requires $\binom{d^{2}}{2}=36$ settings for Alice when the number of Bob’s setting is same as 9 [20].
V Conclusion
We presented a method to construct correlation function for which the quantum maximum is always obtained in a simple analytic form from the generalized EBI correlation. Bell inequality for two-qutrit system derived in our method shows larger violation compared to known Bell inequalities of similar classes, including original EBI. Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI cryptographic protocols based on it.
Acknowledgements.
This research was supported by Korea Institute of Science and Technology Information(KISTI).(No. (KISTI)K24L4M1C2) and the National Research Council of Science & Technology(NST) grant by the Korea government (MSIT) (No. CAP22053-000). J. R. acknowledges support from the National Research Foundation (NRF) of Korea grant funded by the Korea Government (Grant No. NRF-2022M3K2A1083890).
Appendix A Symmetries on the coefficient matrices and optimal measurements
We present the coefficient matrix for $S_{3}$ and MUBs and SICs that gives the optimal violation. Although the symmetries in $S_{3}$ can be expressed with an analytic equation in the main text, the explicit form of the coefficient matrix, MUBs and SICs would show the symmetries in the forms that can be seen at a glance. The coefficient matrices are given as $8\times 9$ matrices satisfying real-valuedness condition, $F_{2}=F_{1}^{*}$ as follows
$F_{1}:=\frac{1}{2}$ $\begin{bmatrix}1&1&1&\omega^{2}&\omega^{2}&\omega^{2}&\omega&\omega&\omega\\ 1&1&1&\omega&\omega&\omega&\omega^{2}&\omega^{2}&\omega^{2}\\ \lambda&\nu&\mu&\lambda&\nu&\mu&\lambda&\nu&\mu\\ \nu&\mu&\lambda&\mu&\lambda&\nu&\lambda&\nu&\mu\\ \mu&\lambda&\nu&\nu&\mu&\lambda&\lambda&\nu&\mu\\ \lambda&\mu&\nu&\lambda&\mu&\nu&\lambda&\mu&\nu\\ \mu&\nu&\lambda&\lambda&\mu&\nu&\nu&\lambda&\mu\\ \nu&\lambda&\mu&\lambda&\mu&\nu&\mu&\nu&\lambda\end{bmatrix}$
when $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. One can find the first column generates the others following, $f_{dp+q,dr+s}^{1}=\omega^{-(ps+qr)}f_{3p+q,0}^{1}$.
[…]
$\mbox{SIC}_{1}$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}0&1&1&0&\omega^{2}&\omega&0&\omega&\omega^{2}\\ 1&0&1&\omega&0&\omega^{2}&\omega^{2}&0&\omega\\ 1&1&0&\omega^{2}&\omega&0&\omega&\omega^{2}&0\end{bmatrix}\,$
$\mbox{SIC}_{2}$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}0&-1&1&0&-\omega^{2}&\omega&0&-\omega&\omega^{2}\\ 1&0&-1&\omega&0&-\omega^{2}&\omega^{2}&0&-\omega\\ -1&1&0&-\omega^{2}&\omega&0&-\omega&\omega^{2}&0\end{bmatrix}$
where each column of $\mbox{SIC}_{1}$ and $\mbox{SIC}_{2}$ defines SICs generated from $\left|\varphi_{1}\right>$ and $\left|\varphi_{2}\right>$ respectively. On the other side, the eigenstates of Alice’s optimal observables $A_{1}=Z$, $A_{3}=X$, $A_{4}=XZ$, $A_{5}=XZ^{2}$ respectively define the complete mutually unbiased bases,
$\mbox{MUB}_{1}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix},\,\mbox{MUB}_{2}=\frac{1}{\sqrt{3}}\begin{bmatrix}1&\omega^{2}&\omega\\ 1&\omega&\omega^{2}\\ 1&1&1\end{bmatrix}$
$\mbox{MUB}_{3}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega^{2}&\omega&1\\ \omega^{2}&1&\omega\\ 1&1&1\end{bmatrix},\,\mbox{MUB}_{4}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega&1&\omega^{2}\\ \omega&\omega^{2}&1\\ 1&1&1\end{bmatrix}.$
We note that the eigenvectors of other settings $A_{2},A_{6},A_{7},A_{8}$ also define the above four MUBs. It can be easily shown from the fact that they are unitary and up to phase equivalent to $A_{1}^{2},A_{3}^{2},A_{5}^{2},A_{4}^{2}$ respectively.
Appendix B Derivation of the quantum bound
Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is to be presented here. The SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as
$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}$ (13)
by substituting $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$. The second term of the RHS of (13) is reduced to $d^{2}(d-1)/2\mathds{1}$ from the unitarity of $B_{y}^{n}$. The first term of the RHS is derived as the desirable form by substituting $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$ into $\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$, as follows,
$\sum_{n,y}\left(\sum_{a}f_{a,y}^{n}A_{a}^{n}\right)^{\dagger}\sum_{x}f_{x,y}^{n}A_{x}^{n}$
$=\sum_{n,a,x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
$=\sum_{n,x,y}|f_{x,y}^{n}|^{2}\mathds{1}+\sum_{n}\sum_{a\neq x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
$=\sum_{n}\lVert F_{n}\rVert^{2}\mathds{1}+\sum_{a\neq x}(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}(A_{a}^{n})^{\dagger}A_{x}^{n}$ (14)
where the Frobeniuns norm $\lVert F_{n}\rVert:=(\sum_{x,y}|f_{x,y}^{n}|^{2})^{1/2}$. Consequently, the upper bound in (13) is reduced to
$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (15)
under conditions $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all $n,a\neq x$.